# Conditional score-based diffusion models for solving inverse problems in mechanics

### v2

@@ 3.4 Design choices for the score network
The design choices for the score networks we use in this work follow suggestions made by Song and Ermon [36]. Primarily, the score network treats $\sigma_{j}$ unlike an usual input such as ${\mathbf{x}}$ or ${\mathbf{y}}$. Rather, $\sigma_{j}$ is used to scale the outputs from the network i.e., $\sigma_{j}$ divides the outcome from the last layer of the score network. Moreover, we treat ${\mathbf{x}}$ and ${\mathbf{y}}$ as images as they are discretized forms of continuous field quantities: ${\mathbf{x}}$ is the spatially varying material constitutive parameter over the specimen discretized over an appropriate Cartesian grid, and ${\mathbf{y}}$ is the measured mechanical response of the specimen over the same Cartesian grid. Therefore, we use convolution neural network type architectures to construct the score functions with ${\mathbf{x}}$ and ${\mathbf{y}}$ as separate channels. More specifically, we use RefineNets [55] to construct the score networks in this work. If there are multiple measurement types available or the measurement type itself has more than one component, we simply add additional channels. For instance, we treat vertical and horizontal displacement measurements as separate channels. Similarly, in the case of a time-dependent problem, if vertical displacement measurements are made at different time instants, the instantaneous measurements are concatenated as additional channels.
@@ 3.5 Workflow
In this section, we describe the workflow for solving an inverse problem using conditional score-based diffusion models. The basic workflow can be divided into three steps; see Fig. 1.
Figure 1: Schematic diagram of the proposed framework for solving inverse problems in mechanics using conditional diffusion models. Step 1 involves creating the dataset by sampling realizations of $({\mathbf{X}},{\mathbf{Y}})$ from the joint distribution $\mathrm{p}_{{\mathbf{X}}{\mathbf{Y}}}$, which requires multiple evaluations of the forward model $\mathcal{F}$. Step 2 involves the training of the score network, which progressively learns to predict the noise in realizations of $\tilde{{\mathbf{X}}}$ conditioned on ${\mathbf{Y}}$ for all levels of noise $\sigma$. Finally, Annealed Langevin dynamics is used to sample from the target posterior distribution in Step 3
@@ 3.5.3 Step 3: Generate samples from the posterior using Annealed Langevin dynamics
In this step, we sample from the target posterior via the Annealed Langevin dynamics algorithm shown in Algorithm 1. First, we specify the number of Langevin steps $T$ and step size $\epsilon$, and use the trained score network $\mathbf{s}(\cdot,\cdot,\cdot;\bm{\theta}^{\prime})$ to generate new realizations of ${\mathbf{X}}$ for a given measurement $\hat{{\mathbf{y}}}$. Algorithm 1 is parallelizable and multiple realizations can be generated together. We use the generated realizations to estimate the posterior mean and posterior standard deviation of ${\mathbf{X}}$, with the former serving as the reconstruction and the latter providing a measure of uncertainty.
@@
Using a trained score network, we can sample the posterior distributions for different realizations of ${\mathbf{Y}}$ without re-training the score network. This makes the proposed approach particularly suitable for high throughput applications requiring repeated solving of the same inverse problem.
@@ 4 Results
In this section, we use conditional score-based diffusion models to solve various large-scale inverse problems in mechanics where the goal is to infer the Young’s modulus (or a related constitutive parameter) field of a material specimen subject to external loading. We consider the following problems:
@@
    Section 4.1.2 — Time-harmonic elastography problem of inferring the shear modulus field of regions surrounding the human optic nerve head from complex measurements of a wave field as the specimen is subject to harmonic excitation.
@@
    Section 4.2.3 — Optical coherence elastography of tumor spheroids (clusters of cells within a hydrogel) wherein the Young’s modulus field of the specimen is inferred from the phase difference between optical coherence microscopy scans as the specimen undergoes compression. This example also involves experimental data and its details can be found in Foo et al. [58].
@@
    Synthetic and real measurements — In Section 4.1, we consider two inverse problems with synthetic measurements. These problems provide the opportunity to validate the proposed method and, for the quasi-static elastography inverse problem in Section 4.1.1 with a relatively simple latent prior distribution, compare it against Monte Carlo inference. In Section 4.2, we apply conditional score-based diffusion models to solve inverse problems with measurements coming from physical experiments. However, note that even in such cases, we train the score network using synthetic data generated from the joint distribution that we model.
@@
We implement the proposed method on PyTorch [59]. Before training the score network, we min-max normalize the training dataset to the range $[0,1]$. We perform this normalization independently for the realizations of ${\mathbf{X}}$ and ${\mathbf{Y}}$ in the training dataset. For the synthetic problems in Section 4.1, we present all results in normalized units. For the inverse problems where we have experimental measurements, we present all results in physical units after inverting the normalization. Additionally, for the tumor spheroid application in Section 4.2.3, we take the logarithm of the Young’s modulus field relative to the Young’s modulus of the surrounding media (hydrogel) before normalizing the realizations of ${\mathbf{X}}$ in the training dataset. This helps improve the scaling of the training data because the Young’s modulus may vary by many orders of magnitude between the cells within a tumor spheroid and the surrounding media. We list the training and sampling hyper-parameters for the various inverse problems in Table 2.
Table 2: Training and sampling hyper-parameters for the conditional diffusion models
@@
Step size $\epsilon$ 5.7$\times$10--6 5.7$\times$10--6 5.7$\times$10--6 5.7$\times$10--6 5.7$\times$10--6
@@ 4.1 Experiments on synthetic data
In this section, we apply the conditional score-based diffusion models (cSDMs) to two inverse problems with synthetic data. These examples help validate the proposed approach by comparing the solutions with Monte Carlo simulation-based inference.
@@ 4.1.1 Quasi-static elastography
The first synthetic example concerns quasi-static elastography, a medical imaging technique in which tissues undergoing deformations under an external load are imaged using ultrasound [23, 60]. The ultrasound images are used to determine the displacements of the heterogeneous specimen, and the goal is to infer the spatially varying shear modulus. This study features a tissue phantom embedded with a stiffer circular inclusion within a softer surrounding matrix, such as the realizations shown in the first row of Fig. 2. In elastography applications, ultrasound can accurately measure the displacement component along the transducer axis, which in this case is the vertical direction. Thus, the inverse problem comprises inferring the spatial distribution of the shear modulus from noisy measurements of vertical displacements of the phantom. The mechanics is governed by the linear elastic model Eq. 22 and 21. Moreover, we assume that the specimen is incompressible and in plane stress condition.
@@
In this example, the dimensions of the specimen is 1$\times$1 cm2. To simulate compression, we fix the bottom edge and impose a downwards vertical displacement of 0.01 cm on the top edge. The bottom edge is traction-free in the horizontal direction, whereas the left and right edges are kept traction-free in both directions. Additionally, we pin the bottom left corner to prevent rigid body motion. A realization of $\bm{X}$ corresponds to the spatial distribution map of the shear modulus of the specimen discretized over a 56$\times$56 Cartesian grid. We formulate a parametric prior that represents a uniformly stiff circular inclusion set against a homogeneous background: the coordinates of the inclusion’s center are uniformly distributed over the domain such that the inclusion fits within the boundaries. We intentionally work with a simple prior, consisting of only two random variables, to facilitate Monte Carlo simulations (MCS). See Table 3 for details regarding the prior distribution for this example. We fix the inclusion’s radius to be 0.12 cm and choose the shear modulus of the inclusion and the background to be 1.5 and 0.1 kPa, respectively.
@@
We also estimate the posterior statistics using Monte Carlo simulation (MCS), which we treat as the true statistics, to validate the results obtained using the cSDMs. We use a sample size equal to 500,000 for the Monte Carlo simulations, which is significantly more than the size of the training dataset. Note that the measurement noise is additive in this case and the likelihood function can be used to obtain the importance weights, which we subsequently self-normalize. We compare the posterior statistics estimated using cSDM and MCS on two test samples that were not part of the training dataset; Fig. 3 shows the results. In Fig. 3(a) and (b), column 1 shows the test phantom while column 2 shows the corresponding measurement for three different noise levels. Note that the ground truth is not readily discernible through visual inspection of the measurements in column 2 of Fig. 3.
Refer to captionRefer to caption(a) Test sample 1(b) Test sample 2
Figure 3: Posterior statistics – pixel-wise posterior mean and standard deviation – estimated using the conditional diffusion models (cSDMs) and Monte Carlo simulation (MCS) for the synthetic quasi-static elastography example on two test samples. In subfigures (a) and (b), the three rows show the results corresponding to three different levels of measurement noise $\sigma_{\eta}$ but for the same ground truth. All values are normalized to $[0,1]$ scale
@@
Fig. 3 also shows the pixel-wise posterior mean and standard deviation estimated using the cSDM and Monte Carlo simulations. We generate 10000, 4000, and 1000 posterior realizations to compute these statistics using the trained score network for $\sigma_{\eta}=$10%, 5%, and 2.5%, respectively. These sample sizes were roughly close to the effective sample size of the Monte Carlo samples for the three cases. Fig. 3 shows that the cSDM mean successfully locates the inclusion at all measurement noise levels, similar to MCS. Furthermore, the posterior statistics estimated using cSDM are qualitatively similar to those obtained using MCS, with the root mean squared errors (RMSE) listed in Table 4. Both these methods predict larger uncertainty around the edges of the inclusion, and the uncertainty predictably grows with the level of measurement noise. For the lowest level of measurement noise, the reconstructed inclusion in the posterior mean is sharp as there is very little uncertainty, both in spread and magnitude, surrounding the inclusion. The cSDM yields results similar to Monte Carlo inference for other test phantoms as well. Table 5 tabulates the RMSE between the posterior statistics estimated using cSDM and MCS across ten test phantoms. The consistently low average RMSE values suggest that the proposed method is accurate and robust to the amount of measurement noise.
@@ 4.1.2 Time-harmonic elastography
In this example, we consider the inverse Helmholtz problem, which arises in elastography applications wherein the heterogeneous mechanical properties of a specimen must be recovered from measurements corresponding to time-harmonic excitation. More specifically, in this example, the shear modulus of the tissue surrounding a typical human optic nerve head (ONH) is to be inferred from measurements of propagating shear waves. Fig. 4 shows a typical realization from the prior and labels the various parts surrounding the ONH. We adapt this example from a previous study by Ray et al. [45].
Figure 4: A realization of the optic nerve head (ONH) sampled from the parametric prior distribution showing the (1) peripapillary sclera, (2) lamina cribrosa, (3) pia matter, (4) optic nerve, and (5) retina. This figure also appears in [45]
@@
Table 6: Ranges of the dimensions for components of the optic nerve head geometry resulting from the prior distribution. These ranges result from combining multiple random variables in some cases
@@
Table 7: Random variables comprising the parametric prior for the optic nerve head (ONH). Note, $\mathcal{N}(\varpi,\varsigma^{2})$ denotes a normal random variable with mean $\varpi$ and variance $\varsigma^{2}$
@@
Figure 5: Five typical realizations of ${\mathbf{X}}$ and ${\mathbf{Y}}$ sampled from the joint distribution that form the training dataset for the synthetic time-harmonic elastography example. Note that ${\mathbf{Y}}$ comprises of ${\mathbf{u}}_{\mathrm{R}}$ and ${\mathbf{u}}_{\mathrm{I}}$ in this case. The first row shows the spatial distribution of the shear modulus field around the optic nerve head, the second and third rows show the corresponding noisy measurements ${\mathbf{u}}_{\mathrm{R}}$ and ${\mathbf{u}}_{\mathrm{I}}$, respectively. All values are normalized to $[0,1]$
We sample the posterior distribution for two test cases that were not part of the training dataset using the trained score network and Annealed Langevin dynamics. For these cases, the first column in Fig. 6 shows the ground truth shear modulus distribution. The second and third columns in Fig. 6 show the corresponding noisy measurements. Note that the shear modulus distribution and the corresponding measurements are all plotted in the normalized $[0,1]$ scale. We sample the target posterior distribution for both cases using Annealed Langevin dynamics and generate 3,000 realizations for each case; Table 2 lists the sampling hyper-parameters for this problem. We use the generated samples to estimate the posterior pixel-wise mean and standard deviation, which we plot in the fourth and fifth columns of Fig. 6, respectively. The last column in Fig. 6 shows the absolute pixel-wise error between the posterior mean estimated using the conditional diffusion model and the corresponding ground truth. The RMSE between the posterior mean and the ground truth are 0.021 and 0.032 for the two test cases, which shows that the posterior mean provides a good reconstruction for the ground truth. We observe relatively larger uncertainty in regions of sharp change in the elastic modulus field, which also coincides with areas of larger reconstruction errors.
Figure 6: Posterior statistics estimated using the conditional diffusion model for the inverse Helmholtz problem with synthetic data. Note, the ground truth distribution of the shear modulus, corresponding measurements and posterior statistics are all plotted in the normalized $[0,1]$ scale

### v3

@@ 3.4 Design choices for the score network
The design choices for the score networks we use in this work follow suggestions made by Song and Ermon [37]. Primarily, the score network treats $\sigma_{j}$ unlike an usual input such as ${\mathbf{x}}$ or ${\mathbf{y}}$. Rather, $\sigma_{j}$ is used to scale the outputs from the network i.e., $\sigma_{j}$ divides the output from the last layer of the score network. Moreover, we treat ${\mathbf{x}}$ and ${\mathbf{y}}$ as images as they are discretized forms of continuous field quantities: ${\mathbf{x}}$ is the spatially varying material constitutive parameter over the specimen discretized over an appropriate Cartesian grid, and ${\mathbf{y}}$ is the measured mechanical response of the specimen over the same Cartesian grid. Therefore, we use convolution neural network type architectures to construct the score functions with ${\mathbf{x}}$ and ${\mathbf{y}}$ as separate channels. More specifically, we use RefineNets [57] to construct the score networks in this work. If there are multiple measurement types available or the measurement type itself has more than one component, we simply add additional channels. For instance, we treat vertical and horizontal displacement measurements as separate channels. Similarly, in the case of a time-dependent problem, if vertical displacement measurements are made at different time instants, the instantaneous measurements are concatenated as additional channels.
@@ 3.5 Workflow
In this section, we describe the workflow for solving an inverse problem using cSDMs. The basic workflow can be divided into three steps; see Fig. 1.
Figure 1: Schematic diagram of the proposed framework for solving inverse problems in mechanics using cSDMs. Step 1 involves creating the dataset by sampling realizations of $({\mathbf{X}},{\mathbf{Y}})$ from the joint distribution $\mathrm{p}_{{\mathbf{X}}{\mathbf{Y}}}$, which requires multiple evaluations of the forward model $\mathcal{F}$. Step 2 involves the training of the score network, which progressively learns to predict the noise in realizations of $\tilde{{\mathbf{X}}}$ conditioned on ${\mathbf{Y}}$ for all levels of noise $\sigma$. Finally, Annealed Langevin dynamics is used to sample from the target posterior distribution in Step 3
@@ 3.5.3 Step 3: Generate samples from the posterior using Annealed Langevin dynamics
In this step, we sample from the target posterior via the Annealed Langevin dynamics algorithm shown in Algorithm 1. First, we specify the number of Langevin steps $T$ and step size parameter $\epsilon$, and use the trained score network $\mathbf{s}(\cdot,\cdot,\cdot;\bm{\theta}^{\prime})$ to generate new realizations of ${\mathbf{X}}$ for a given measurement $\hat{{\mathbf{y}}}$. Algorithm 1 is parallelizable and multiple realizations can be generated together. In practice, and for the numerical experiments in this paper, posterior realizations are generated in batches and, hence, in parallel. The batch size for sampling is not the same as, and often greater than, the batch size for training. The dimensionality of the inverse problem at hand and the memory capacity of the compute resource, CPU or GPU, dictates the sampling batch size. We use the generated realizations to estimate the posterior mean and posterior standard deviation of ${\mathbf{X}}$, with the former serving as the reconstruction and the latter providing a measure of uncertainty.
@@
cSDMs only require samples from the joint distribution $\mathrm{p}_{{\mathbf{X}}{\mathbf{Y}}}$ between ${\mathbf{X}}$ and ${\mathbf{Y}}$ to train the conditional score network. Hence, the proposed approach is effectively a likelihood-free simulation-based inference approach [58]. It is well known that the sampling efficiency of many traditional likelihood-free approaches scales poorly with the dimensionality of the inverse problem [58]. In contrast, the proposed approach scales with the dimensionality of the inverse problem at hand due to the neural network-based score function approximation. This scalability stems from the adeptness of the neural network in handling high-dimensional data and modeling complex patterns between them. Moreover, the trained conditional score network approximates the posterior’s score function $\nabla_{{\mathbf{x}}}\log\mathrm{p}_{{\mathbf{X}}|{\mathbf{Y}}}\!\left({{\mathbf{x}}|\hat{{\mathbf{y}}}}\right)$ for all realizations of $\hat{{\mathbf{y}}}$. Therefore, using a trained score network, we can sample the posterior distributions for different realizations of ${\mathbf{Y}}$ without re-training the score network. This helps ‘amortize’ the inference cost over ${\mathbf{Y}}$ [59] and makes the proposed approach particularly suitable for high throughput applications requiring repeated solving of the same inverse problem. A comparative study of cSDMs with other simulation-based and likelihood-free inference methods is beyond the scope of the current work.
@@ 4 Results
In this section, we use cSDMs to solve various large-scale inverse problems in mechanics where the goal is to infer the Young’s modulus (or a related constitutive parameter) field of a material specimen subject to external loading. We consider the following problems:
@@
    Section 4.1.2 — Time-harmonic elastography problem of inferring the shear modulus field of regions surrounding the human optic nerve head (ONH) from complex measurements of a wave field as the specimen is subject to harmonic excitation.
@@
    Section 4.2.3 — Optical coherence elastography (OCE) of tumor spheroids (clusters of cells within a hydrogel) wherein the Young’s modulus field of the specimen is inferred from the phase difference between optical coherence microscopy scans as the specimen undergoes compression. This example also involves experimental data and its details can be found in Foo et al. [62].
@@
    Synthetic and real measurements — In Section 4.1, we consider two inverse problems with synthetic measurements. In particular, the quasi-static elastography inverse problem in Section 4.1.1 has a relatively simple latent prior distribution amenable to Monte Carlo simulation (MCS), which provides an opportunity to validate the proposed method. In Section 4.2, we apply cSDMs to solve inverse problems with measurements coming from physical experiments. However, note that even in such cases, we train the score network using synthetic data generated from the joint distribution that we model.
@@
We implement the proposed method on PyTorch [63]. Before training the score network, we min-max normalize the training dataset to the range $[0,1]$. We perform this normalization independently for the realizations of ${\mathbf{X}}$ and ${\mathbf{Y}}$ in the training dataset. For the synthetic problems in Section 4.1, we present all results in normalized units, meaning all posterior statistics are also estimated in the normalized units. For the inverse problems where we have experimental measurements, we present all results in physical units after inverting the normalization. Additionally, for the tumor spheroid application in Section 4.2.3, we take the logarithm of the Young’s modulus field relative to the Young’s modulus of the surrounding media (hydrogel) before normalizing the realizations of ${\mathbf{X}}$ in the training dataset. This helps improve the scaling of the training data because the Young’s modulus may vary by many orders of magnitude between the cells within a tumor spheroid and the surrounding media. We list the training and sampling hyper-parameters for the various inverse problems in Table 2.
Table 2: Training and sampling hyper-parameters for the cSDMs
@@
Step size parameter $\epsilon$ 5.7$\times$10--6 5.7$\times$10--6 5.7$\times$10--6 5.7$\times$10--6 5.7$\times$10--6
@@
Table 3: Approximate computational cost of using cSDMs for Bayesian inference
Inverse problem Step 1$\ast$ Step 2$\dagger$ Step 3$\ddagger$
Wall time Resource Wall time Resource Wall time Resource
Circular inclusion (Section 4.1.1) 11 s 4 $\times$ 12 core Intel Xeon 4116 CPU 43 s NVIDIA A40 GPU 2 mins 2 $\times$ NVIDIA P100 GPU
Optic nerve head (Section 4.1.2) 17 min 8 core Apple M1 Pro 30 s NVIDIA A40 GPU 2 mins NVIDIA A40 GPU
Circular inclusion (Section 4.2.1) 11 s 4 $\times$ 12 core Intel Xeon 4116 CPU 1 min NVIDIA Quadro RTX 8000 3 mins NVIDIA Quadro RTX 8000
Dual-layer phantom (Section 4.2.2) 1.5 hrs 4 $\times$ 4 core Intel i5 7400 CPU 35 s NVIDIA A100 GPU 90 s NVIDIA A100 GPU
Tumor spheroids (Section 4.2.3) 2.5 $\star$ 2 $\times$ 8 core Intel Xeon E5-2690 CPU 2 mins 2 $\times$ NVIDIA A40 GPU 45 mins 2 $\times$ NVIDIA A40 GPU
[…]
Breakdown of computational costs for various steps in the workflow
Table 3 provides an estimate of the wall times (a proxy for the computational cost) for the three steps in the workflow and the resource type used. In general, we make the following remarks. The cost of generating the training dataset depends on the complexity and dimensionality of the problem. The cost of training the score network depends on the batch size and the inverse problem’s dimensionality. The time required for posterior sampling is influenced by the total number of Langevin steps, which equals $L\times T$. This is perhaps the main drawback of using diffusion models compared to other generative models and remains an active research area.
@@ 4.1 Experiments on synthetic data
In this section, we apply cSDMs to two inverse problems with synthetic data. The inverse problem in Section 4.1.1 helps validate the proposed approach by comparing the solutions with MCS.
@@ 4.1.1 Quasi-static elastography
The first synthetic example concerns quasi-static elastography, a medical imaging technique in which tissues undergoing deformations under an external load are imaged using ultrasound [23, 64]. The ultrasound images are used to determine the displacements of the heterogeneous specimen, and the goal is to infer the spatially varying shear modulus. This study features a tissue phantom embedded with a stiffer circular inclusion within a softer surrounding matrix, such as the realizations shown in the first row of Fig. 2. In elastography applications, ultrasound can accurately measure the displacement component along the transducer axis, which in this case is the vertical direction. Thus, the inverse problem comprises inferring the spatial distribution of the shear modulus from noisy measurements of vertical displacements of the phantom. The mechanics is governed by the linear elastic model Eq. 22 and 21. Moreover, we assume that the specimen is incompressible and in plane stress condition following previous studies [42, 47].
@@
In this example, the dimensions of the specimen is 1$\times$1 cm2. To simulate compression, we fix the bottom edge and impose a downwards vertical displacement of 0.01 cm on the top edge. The bottom edge is traction-free in the horizontal direction, whereas the left and right edges are kept traction-free in both directions. Additionally, we pin the bottom left corner to prevent rigid body motion. A realization of $\bm{X}$ corresponds to the spatial distribution map of the shear modulus of the specimen discretized over a 56$\times$56 Cartesian grid. We formulate a parametric prior that represents a uniformly stiff circular inclusion set against a homogeneous background: the coordinates of the inclusion’s center are uniformly distributed over the domain such that the inclusion fits within the boundaries. We intentionally work with a simple prior, consisting of only two random variables, to facilitate MCS. See Table 4 for details regarding the prior distribution for this example. We fix the inclusion’s radius to be 0.12 cm and choose the shear modulus of the inclusion and the background to be 1.5 and 0.1 kPa, respectively.
@@
We also estimate the posterior statistics using MCS, which we treat as the true statistics, to validate the results obtained using the cSDMs. We use a sample size equal to 500,000 for the MCS, which is significantly more than the size of the training dataset. Note that the measurement noise is additive in this case and the likelihood function can be used to obtain the importance weights, which we subsequently self-normalize. See Appendix B for details of the MCS procedure. We compare the posterior statistics estimated using cSDM and MCS on two test samples that were not part of the training dataset; Fig. 3 shows the results. In Fig. 3(a) and (b), column 1 shows the test phantom while column 2 shows the corresponding measurement for three different noise levels. Note that the ground truth is not readily discernible through visual inspection of the measurements in column 2 of Fig. 3.
@@
Refer to captionRefer to caption(a) Test sample 1(b) Test sample 2
Figure 3: Posterior statistics – pixel-wise posterior mean and standard deviation – estimated using cSDMs and MCS for the synthetic quasi-static elastography example on two test samples. In subfigures (a) and (b), the three rows show the results corresponding to three different levels of measurement noise $\sigma_{\eta}$ but for the same ground truth. All values are normalized to $[0,1]$ scale
@@
Fig. 3 also shows the pixel-wise posterior mean and standard deviation estimated using the cSDM and MCS. We generate 10000, 4000, and 1000 posterior realizations to compute these statistics using the trained score network for $\sigma_{\eta}=$10%, 5%, and 2.5%, respectively. These sample sizes were roughly close to the effective sample size of the MC samples for the three cases. Fig. 3 shows that the posterior mean estimated using cSDM successfully locates the inclusion at all measurement noise levels, similar to MCS. Furthermore, the posterior statistics estimated using cSDM are qualitatively similar to those obtained using MCS, with the root mean squared errors (RMSE) listed in Table 5. Both these methods predict larger uncertainty around the edges of the inclusion, and the uncertainty predictably grows with the level of measurement noise. For the lowest level of measurement noise, the reconstructed inclusion in the posterior mean is sharp as there is very little uncertainty, both in spread and magnitude, surrounding the inclusion. The cSDM yields results similar to MCS for other test phantoms as well. Table 6 tabulates the RMSE between the posterior statistics estimated using cSDM and MCS across ten test phantoms. The consistently low average RMSE values suggest that the proposed method is accurate and robust to the amount of measurement noise.
Effect of model misspecfication
To understand the effects of a misspecified model for $\mathrm{p}_{{\mathbf{Y}}|{\mathbf{X}}}$, we perform inference for the case when the measurement noise’s standard deviation $\sigma_{\eta}=0.025$ and $0.1$ with the score network trained assuming $\sigma_{\eta}=0.05$. For test sample 1, Fig. 4 shows the posterior statistics estimated using the cSDM and compares it against the MCS under the misspecified noise model. Comparing Tables 7 and 5, the RMSE in the posterior statistics estimated using the cSDM increases when the noise model is misspecified. In particular, Fig. 4 and Table 7 show that the pixel-wise posterior mean is particularly biased when the noise in the measurement has standard deviation $\sigma_{\eta}=0.1$ but inference is performed assuming $\sigma_{\eta}=0.05$. These results highlight the need for studying the effects of model misspecification on the performance of cSDMs.
Figure 4: Posterior statistics – pixel-wise posterior mean and standard deviation – for test sample 1 estimated using cSDMs trained assuming measurement noise with standard deviation $\sigma_{\eta}=0.05$. The reference MC statistics are also estimated with the misspecified measurement noise model. All values are normalized to $[0,1]$ scale
Table 7: Root mean squared error (RMSE) between the pixel-wise posterior statistics estimated using cSDM and MCS for misspecified values of $\sigma_{\eta}$
Measurement noise ($\sigma_{\eta}$) Mean Std. Dev.
10.0% 0.069 0.038
 2.5% 0.024 0.037
@@ 4.1.2 Time-harmonic elastography
In this example, we consider the inverse Helmholtz problem, which arises in elastography applications wherein the heterogeneous mechanical properties of a specimen must be recovered from measurements corresponding to time-harmonic excitation. More specifically, in this example, the shear modulus of the tissue surrounding a typical human ONH is to be inferred from measurements of propagating shear waves. Fig. 5 shows a typical realization from the prior and labels the various parts surrounding the ONH. We adapt this example from a previous study by Ray et al. [47].
@@
Figure 5: A realization of the ONH sampled from the parametric prior distribution showing the (1) peripapillary sclera, (2) lamina cribrosa, (3) pia matter, (4) optic nerve, and (5) retina. This figure also appears in [47]
Table 8: Ranges of the dimensions for components of the ONH geometry resulting from the prior distribution. These ranges result from combining multiple random variables in some cases
@@
Table 9: Random variables comprising the parametric prior for the ONH. Note, $\mathcal{N}(\varpi,\varsigma^{2})$ denotes a normal random variable with mean $\varpi$ and variance $\varsigma^{2}$
@@
Figure 6: Five typical realizations of ${\mathbf{X}}$ and ${\mathbf{Y}}$ sampled from the joint distribution that form the training dataset for the synthetic time-harmonic elastography example. Note that ${\mathbf{Y}}$ comprises of ${\mathbf{u}}_{\mathrm{R}}$ and ${\mathbf{u}}_{\mathrm{I}}$ in this case. The first row shows the spatial distribution of the shear modulus field around the ONH, the second and third rows show the corresponding noisy measurements ${\mathbf{u}}_{\mathrm{R}}$ and ${\mathbf{u}}_{\mathrm{I}}$, respectively. All values are normalized to $[0,1]$
Figure 7: Posterior statistics estimated using the cSDMs for the inverse Helmholtz problem with synthetic data. Note, the ground truth distribution of the shear modulus, corresponding measurements and posterior statistics are all plotted in the normalized $[0,1]$ scale
We sample the posterior distribution for two test cases that were not part of the training dataset using the trained score network and Annealed Langevin dynamics. For these cases, the first column in Fig. 7 shows the ground truth shear modulus distribution. The second and third columns in Fig. 7 show the corresponding noisy measurements. Note that the shear modulus distribution and the corresponding measurements are all plotted in the normalized $[0,1]$ scale. We sample the target posterior distribution for both cases using Annealed Langevin dynamics and generate 3,000 realizations for each case; Table 2 lists the sampling hyper-parameters for this problem. We use the generated samples to estimate the posterior pixel-wise mean and standard deviation, which we plot in the fourth and fifth columns of Fig. 7, respectively. The last column in Fig. 7 shows the absolute pixel-wise error between the posterior mean estimated using the cSDM and the corresponding ground truth. The RMSE between the posterior mean and the ground truth are 0.021 and 0.032 for the two test cases, which shows that the posterior mean provides a good reconstruction for the ground truth. We observe relatively larger uncertainty in regions of sharp change in the elastic modulus field, which also coincides with areas of larger reconstruction errors.